Fermi National Laboratory
 

Operation Summary for July 11, 2011 - July 13, 2011

 

Monday July 11

The day shift began with NuMI off due to target replacement, with the Recycler stash gone due to a lightning strike, with Operations monitoring the Antiproton Source (Pbar) stack (45.1E10), with the Booster off due to a West Anode Power Supply (WAPS) trip, with no beam going to the Main Injector (MI), or to the MiniBooNE or Meson FTBFT-1015 experiments, with Linac being the only accelerator operating normally, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store.

During the day shift, Booster personnel reported that they brought WAPS back on line at 8:19 AM. Operations resumed sending beam to MiniBooNE. Operators reported that the Tevatron (TEV) ramp was lost due to a lightning strike. Operations established beam to T-1015 at 8:50 AM. A Controls person reported at 8:57 AM that a lightning strike took out the fiber clock repeater upstairs in MI-31 and downstairs the clock fan-out chassis. Operators had D0 and CDF prepared for access at 9:32 AM. Water technicians took the TEV sectors B4 and F2 heat exchangers off line for maintenance. MI personnel began working on their coalescing problems. A TEV Kicker person completed his work on the E48 kicker at 11:13 AM. D0 and CDF were out of access by 11:57 AM. At 12:11 PM, operators tried to transfer antiprotons from Pbar to the Recycler, but the efficiency was poor. Cryo system techs reported at 1:29 PM that they had completed work on the TEV sectors A4 and B2 wet engines. Electron Cooling personnel reported finding a bad amplifier in the Pelletron pit that was inhibiting transfers. Operators put the TEV through a dry squeeze at 3:26 PM.

During the evening shift, operators put the TEV through a wet squeeze at 4:05 PM. Operators reported at 4:13 PM that the TEV D4 LCW pump had tripped off. Water techs investigated and reported that the pump would not start; operators contacted FESS personnel who turned on the pump, but reported that they thought the heat exchanger was clogged. MI personnel reported 4:53 PM that they inspected the MI D5 longitudinal damper system and could not find anything wrong with it; then the system began to work normally. Operations put the TEV into shot setup at 5:41 PM and established store 8880 with an average initial luminosity (AIL) of 162.82E30. Operations disabled beam to T-1015 at 9:01 PM. Operators reported short downtimes due to trips by two Linac quadrupole power supplies (QPS105 and QPS115) caused by an I- Source spark and due to Linac and Booster RF station trips (LRF4 and BRF3).

Tuesday July 12

The midnight shift began with Operations monitoring store 8880 (71.26E30) and the stash (146.7E10), with Pbar stacking (20.5E10), and with MiniBooNE taking beam.

During the midnight shift, at 12:38 AM, the TEV sector E3 cryo system and wet engine came into alarm. Operators aborted the store and dumped the ramp at 12:51 AM. Operators investigated and reported that the wet engine had lost its belt. Cryo system personnel reported at 2:18 AM that the E3 wet engine also required an overhaul, which would take about eight hours. Operations established beam to T-1015 at 4:27 AM. At 7:40 AM, BRF8 tripped off for the first of four times in two minutes. Booster RF personnel investigated.

During the day shift, a MiniBooNE person reported at 9:47 AM that the temperature was climbing in the MI-12 service building; the Duty Mechanic investigated and found one of the air conditioners had frozen, which they thawed out. Water techs worked on more heat exchangers during the day. MI personnel conducted a successful coalescing check at 11:06 AM. At 11:36 AM, the MI Operations Specialist reported finding all the MI transverse dampers off; everything reset. Operations received permission to send beam to MTA through its nitrogen gas filled RF cell. Cryo system techs reported at 12:28 PM that they had completed their overhaul of the E3 wet engine. Operators prepared the TEV Transfer Hall for an access at 1:24 PM to check for air leaks; operators had reported problems closing some valves. An EE Support person spent a half hour working on a MI Lambertson power supply (I:LAM10), which had been tripping off. Operators reported at 2:08 PM that Booster personnel were taking down the Gradient Magnet Power Supply (GMPS) to check for LCW leaks; this work halted beam to MiniBooNE for about an hour. At 3:20 PM, I:LAM10 tripped off again; an EE Support person investigated. Operations began sending beam to MTA at 3:49 PM. A TEV Mechanical tech reported at 3:42 PM, from the Transfer Hall, that a copper air line buried under cables was leaking; he will run a new line.

During the evening shift, EE Support personnel reported at 4:23 PM that he had brought I:LAM10 back online after adjusting a phase unbalance relay. Operations disabled beam to MTA at 4:35 PM; the experimenters reported that the first results were close to expectations. Operators reported that MI-12 temperatures were rising; the Duty Mechanic investigated and reported that all the air conditioners were on. At 6:54 PM, the Pelletron tripped off; operators brought it back online. Operations disabled beam to T-1015 at 9:47 PM; this was their last day of beam for this run. They will be back in December. Operations put the TEV into shot setup at 9:47 PM and established store 8884 at 10:15 PM with an AIL of 394.73E30. Operators reported short downtimes due to a Linac quadrupole power supply (QPS115) trip and a LRF1 trip and a vacuum burst in the MI-8 line.

Wednesday July 13

The midnight shift began with Operations monitoring store 8884 (251.75E30) and the stash (174.4E10), with Pbar stacking (11.7E10), and with MiniBooNE taking beam.

During the midnight shift, operators reported at 3:07 AM that beam positions in the Linac were moving. Operators tried various tuning changes and adjustments over the next three hours, but nothing helped the situation; they returned everything to the original settings. At 7:46 AM, operators reported a reduced efficiency of antiproton transfers; operators tuned and this helped some.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow a MI access, after the next store goes in, to look for a LCW leak. Booster will use this time to access and change out a power amplifier. Linac personnel will investigate the cause of the I- Source sparking.
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/13/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory